OpenStudy (jcarangui1):

Can someone help to find the value of tan (4pi/3) without using a calculator ?

geerky42 (geerky42):

Unit Circle Chart is your friend.

geerky42 (geerky42):



OpenStudy (anonymous):

\[\tan \frac{ 4 \pi }{ 3 }=\tan \left( \pi+\frac{ \pi }{ 3 } \right)=\tan \frac{ \pi }{ 3 }=\frac{ \sin \frac{ \pi }{ 3 } }{ \cos \frac{ \pi }{ 3 } }\] \[=\frac{ \frac{ \sqrt{3} }{ 2 } }{ \frac{ 1 }{ 2 } }=\sqrt{3}\]

OpenStudy (jdoe0001):

or use your Unit Circle

OpenStudy (triciaal):

@excellency I think you missed "without using a calculator "